OpenStudy (anonymous):

Can somebody help me with a test? I will medal them..its just some of the questions won't go here due to the fact its like a fraction..

OpenStudy (anonymous):

Question 1

OpenStudy (abmon98):

7x=-22-6 x=-4

OpenStudy (anonymous):

Ok thank you Heres question 2: )

OpenStudy (abmon98):

10y=30 y=3

OpenStudy (anonymous):

Heres question 3, theres two more questions after this and ill medal you :]

OpenStudy (abmon98):

(12*3)-2=34 -z=34 z=-34

OpenStudy (anonymous):

OK, thank you! :3 1 more after 4 :]

OpenStudy (lena772):

Posting test questions is cheating and against the CoC

OpenStudy (anonymous):

Oh, sorry. Im still new here D<

OpenStudy (lena772):

it's okay. take a look at the coc here. http://openstudy.com/code-of-conduct & welcome to openstudy!

OpenStudy (anonymous):

Thank you!